# Boundary layers, transport and universal distribution in boundary driven active systems

### v2

@@ Abstract
We discuss analytical results for a run-and-tumble particle (RTP) in one dimension in presence of boundary reservoirs. It exhibits ‘kinetic boundary layers’, nonmonotonous distribution, violation of Fourier’s law, diffusion facilitated current reversal and optimisation on tuning dynamical parameters, and a Seebeck-like effect in the steady state. The spatial and internal degrees of freedom together possess a symmetry, using which we find the eigenspectrum for large systems. The eigenvalues are arranged in two bands which can mix in certain conditions resulting in a crossover in the relaxation. The late time distribution for large systems is obtained analytically; it retains a strong and often dominant ‘active’ contribution in the bulk rendering an effective passive-like description inadequate. A nontrivial ‘Milne length’ also emerges in the dynamics. Finally, a novel universality is proposed in the absorbing boundary problem for dynamics with short-ranged colored noise. Active particles driven by active reservoirs may thus provide a common physical ground for diverse and new nonequilibrium phenomena.
Pritha Dolai1,2,3 and Arghya Das4$\star$
@@ 1 Introduction
Systems with self-propelled or active particles became a paradigm of nonequilibrium processes. Active agents consume energy from environment and convert it into mechanical work to propel themselves. Constant energy consumption drives active systems far from equilibrium [1, 2, 3]. There are numerous real life examples of active systems across scales including bacterial colony, school of fish, flock of birds, janus particles[4, 5], vibrated granular rods and disks [6, 7] and many more. Such systems exhibit a plethora of nontrivial collective phenomena, e.g. pattern formation, motility-induced phase separation and clustering [2, 3, 8, 9, 10], giant number fluctuations [11], casimir effect [12]. They reveal rich nonequilibrium properties even at the single particle level: examples include non-Boltzmann steady-state, accumulation near boundary, shape transition in probability distribution in a trapping potential, a condensation-like transition in the free motion with random activity [13, 14, 15, 16, 17, 18]. Active motion is modelled with a Langevin-like equation with an ‘active noise’, which is typically coloured and violates the fluctuation-response relation, leading to such novel behaviour. In literature there are three widely studied scalar active particle models: run-and-tumble particles (RTP), active Brownian particles (ABP) and active-Ornstein-Uhlenbeck particles (AOUP).
@@
Active particles show interesting behaviour in the presence of drives, thermal noise, obstacles and different boundary conditions. For example, RTPs subject to spatially periodic drive go through a transition from non-ergodic trapped states to moving states [19]. In presence of shear forces in two dimension, underdamped ABPs develop a boundary layer and flow reversal occurs near the wall [20]. Confining boundaries are known to have strong effects on active systems, e.g. boundary accumulation and boundary layer formation [13, 21], long-range effects and system size dependent behaviour [22, 23, 24, 25, 26, 27], etc. Interestingly, the nature of the boundary layer and the pressure exerted by the particles on a wall depends on the shape of the wall [21] as well as on the interactions [22]. Thermal fluctuations, ubiquitous but usually neglected, can have nontrivial effects on active motion e.g. formation of bound states [28], current modulation on a ratchet [29], and triggering phase separation [30, 31].
@@
In this work we focus on boundary-driven processes. Systems connected to particle or energy reservoirs have been studied for centuries, but there is a renewed interest in these processes particularly in the context of low dimensional transport. The canonical picture is given by Brownian motion in a finite system connected to particle baths: the steady state density profile is a linear interpolation of the densities at the boundaries [32] while the current is proportional the density gradient as given by the Fick’s law. However in the presence of interaction, bulk drive, or more than one conserved quantities etc. richer and often nonintuitive behaviour is observed. Models of boundary driven interacting diffusive systems exhibit long ranged correlation [33, 34, 35]; introduction of a bulk drive gives rise to different phases [36, 37]. The thermal transport properties become anomalous and often violate Fick’s law in one dimensional interacting anharmonic chains [38, 39, 40]. Nonlinear bulk density profile and nongradient steady state current emerges in simple noninteracting cases where the particle dynamics is governed by strongly correlated noise, for instance Levy walk connected to particle reservoirs [41, 40]. Violation of the Fourier’s law i.e. a finite thermal conduction with zero temperature gradient was demonstrated for nonlinearly coupled passive systems driven by non-Gaussian white noises [43, 42]. New effects induced by ‘active heat baths’ in a passive harmonic (linearly coupled) chain, like current reversal in the energy transport and boundary features in the steady state, have also been reported recently [44, 45]. In contrast, results for active systems in presence of boundary reservoirs are rare[46], and to our knowledge the effects of active reservoirs are not yet discussed. It is natural to ask what are the steady state, transport properties, dynamical fluctuations and relaxation behaviour of different boundary driven active systems and how do they differ across models and from their passive counterpart. To explore the nonequilibrium physics in this class of processes, we consider an elementary model amenable to simple analysis and exact results.
@@
In this paper we study non-interacting free RTPs in presence of thermal noise in a one-dimensional channel connected to particle reservoirs at both ends maintaining fixed particle densities and magnetisation  [^1] Since the RTPs carry two internal states in $1$-dimension, the magnetisation is the difference of densities of the two species.. The results are surprisingly rich and intriguing. The solution for steady state and large time distribution show the existence of ‘kinetic boundary layers’ [47] characterised by exponentially decaying components in the density profile near the reservoirs. These are distinct from the usual boundary layers reported earlier for confining walls or obstacles in that, here these occur in the presence of particle fluxes accross the boundaries and can correspond to both accumulation and depletion. For zero boundary magnetisation, the steady state current and bulk density profile show a passive-like behaviour while the effect of activity which is notable for large persistence, becomes prominent at the boundary layer. However, the signature of activity appears in the bulk through a finite magnetisation. Remarkably, in the presence of nonzero boundary magnetisations, the density profile can become nonmonotonic, there is current without density gradient, and in particular a diffusion-facilitated current reversal and current maximisation upon tuning the particle parameters can take place. In this connection a Seebeck-like effect is identified, with boundary magnetisations taking the role of temperatures and density differences being analogous to voltage differences.
@@
The time-dependent solution to the boundary value problem is identical to that of the absorbing boundary problem which is relatively less discussed [48, 49, 50]. We note that the boundary condition endows the system with a ‘reflection symmetry’. Using this, the eigenspectrum and time-dependent solutions are obtained analytically for large system-sizes. The eigenvalues of the time evolution operator is arranged in two distinct bands separated by the tumble rate. The relaxation is diffusive for large $L$; as the system size is reduced the bands start overlapping, and for small enough systems the relaxation rate crosses over to an $L$-independent value. Such crossovers had been found earlier for one and two particle systems on a ring[51, 28]. We observe that in each of the eigenstate the ‘active’ contribution appears only as $O(L^{-1})$ correction to the leading ‘passive-like’ term for large $L$; but, contrary to expectations, at large times both the active and passive contributions occur at the same order in the bulk. For higher persistence, the large distance late time distribution and related properties are in fact dominated by the active contribution. Interestingly, although the distribution satisfies the absorbing condition at the boundaries in the presence of thermal noise, a ‘Milne length’ [52, 53, 54] is still identified in the distribution. This length has also been found in the context of survival probability of random flights [55] and athermal RTPs [49]. In the present case it dependends nontrivially on the diffusivity and other dynamical parameters. We found rather non-intuitively that the Milne length has a significant role in the the steady state. A kinetic boundary layer related to the Milne length appears in the time-dependent solution as well and this is argued to be a consequence of thermal diffusion. We further argue that this is related to a net imbalance of the two species of particles in the system, which goes away as soon as the diffusivity vanishes. As a whole, the interplay of thermal noise and active motion causes important ramifications in the dynamics in the presence of boundary driving.
@@
It turns out that many of the said features obtained for RTPs also appear in the cases involving underdamped passive particle, athermal AOUP, and ABP, and are possibly generic to dynamics with coloured noise. In the absorbing boundary problem the emergence of kinetic boundary layers and Milne length were known long back for the underdamped passive motion [54, 56, 57, 58]. In that case the large time density profile is in fact quantitatively similar to that obtained here for RTPs. It prompts us to propose that the late time distribution away from the absorbing wall is independent of the realisation of the coloured noise, provided that the noise correlation is short ranged.
@@
Rest of the paper is organized as follows. In Section 2, we describe the model system of a boundary-driven run-and-tumble particle. The steady-state and the time-dependent solutions are presented in Sections  3 and  4, respectively. A universality in the large time distribution is proposed in Section 5. In Section 6 we summarise our perspective and conclude.
@@ 2 Model: boundary driven RTPs
We consider a run-and-tumble particle (RTP) moving in a finite one-dimensional box of length $L$ subject to a translational noise $\eta(t)$. The box is connected to particle reservoirs at the two ends. We assume that the microscopic details of the interaction of the reservoir with the system at the boundary is not important except that it provides a fixed boundary condition. The equation of motion of the particle is,
@@
where $v$ is the self-propulsion speed, $\sigma(t)$ is the intrinsic orientation (spin) characterized by a telegraphic noise that takes values ±$\pm 1$ and it flips between the two states with rate $\omega$, and $\eta(t)$ is the Gaussian white noise with $\langle\eta(t)\rangle=0,~{}\langle\eta(t)\eta(t^{\prime})\rangle=2D\delta(t-t^{\prime}),~{}D$ being the diffusion constant.
Figure 1: Schematic of the boundary driven active system. Each RTP moves on a one-dimensional line bounded between $x=0$ and $L$. The system is connected to particle reservoirs at both the ends maintaining fixed particle densities $P_{0}$ and $P_{1}$ and the corresponding particle magnetisation $Q_{0}$ and $Q_{1}$ at $x=0,\,L$ respectively.
The run and tumble motion described in terms of $(x(t),\sigma(t))$ is a Markov process. The state of the particle at a time $t$ is given by, $|P(x,t)\rangle=\begin{bmatrix}P_{+}(x,t)\\ P_{-}(x,t)\end{bmatrix}$, where $P_{+}(x,t)$ and $P_{-}(x,t)$ are the probabilities to find a particle at x at time t with instantaneous spin $+1$ and $-1$ respectively. $P_{+}$ and $P_{-}$ satisfy the Master equations,
@@
$P_{0}^{\pm}$ and $P_{1}^{\pm}$ are the probability densities of each species of particles at the two boundaries $x=0$ and $x=L$ respectively. The total probability density is $P(x,t)=P_{+}(x,t)+P_{-}(x,t)$. We also define a quantity $Q(x,t):=\langle\sigma\rangle_{(x,t)}=P_{+}(x,t)-P_{-}(x,t)$ which is analogous to magnetization density. In terms of $P$ and $Q$ the boundary conditions are rewritten as,
@@ 3 Steady state distribution
$\partial_{t}P=D\partial_{x}^{2}P-v\partial_{x}Q=0$ (5)
$\partial_{t}Q=D\partial_{x}^{2}Q-v\partial_{x}P-2\omega Q=0.$ (6)
For arbitrary densities of the two species of particles at each of the boundaries, we need to solve Eqs. (5)-(6) with the boundary conditions given in Eq. (4). The solution for the density and magnetisation profiles and the current $J_{0}=-D\partial_{x}P+v\,Q$ in the steady state are,
@@
where, $\Delta P=P_{1}-P_{0},~{}\Delta Q=Q_{1}-Q_{0}$, $D_{e}=D+\frac{v^{2}}{2\omega}$, $\mu=\frac{\sqrt{2\omega D_{e}}}{D}$, $M=\frac{v}{\mu D}\frac{1-e^{-\mu L}}{1+e^{-\mu L}}$, and $B_{0}=\frac{\Delta P-M(Q_{0}+Q_{1})}{2(M+L\frac{\omega}{v})}+Q_{0}-\Delta Q\,\frac{e^{-\mu L}}{1-e^{-\mu L}}\,$. In the $D=0$ case $\mu$ diverges and the exponential terms vanish. This gives rise to discontinuities in the density and magnetisation profiles at the boundaries and it is smoothened to kinetic boundary layers for nanvanishing diffusion. We list a few implications of the Eqs. (7)-(9).
  • I.
    For $L$ large the current and the density gradient in the bulk ($\mu^{-1}\ll x\ll L$) are proportional to $(L+2\,l_{M})^{-1}$, where $l_{M}=\frac{v^{2}}{2\omega\,\sqrt{2\omega D_{e}}}$ is the Milne length as shown later. This increase in the effective system length suggests that the system in the bulk pertains to boundary conditions at a distance $l_{M}$ beyond the actual boundaries, as in several absorbing boundary problems. This picture however holds only for $Q_{0}=Q_{1}=0$.
  • II.
    The bulk density profile, when extended to the boundaries, violates the boundary condition by an amount proportional to $M$ (or $l_{M}$). For $D>0$ ‘fine-tuned’ kinetic boundary layers representing particle accumulation at one end and depletion at the other are formed which restore the boundary conditions. Further, the density profile may become nonmonotonic near the boundaries.
  • III.
    The current is finite even if $\Delta P=0$: $J_{0}\propto(Q_{0}+Q_{1})/(L+2\,l_{M})$, in violation of Fourier’s law. This suggests that the boundary magnetisations effectively act as drives inducing a constant current in the system in the absence of a density gradient. However this similarity cannot be exact, since in the present case the current gives rise to a linear density profile away from the boundaries, unlike the usual driven passive case in which density remains uniform.
  • IV.
    On the contrary, with suitable boundary magnetisations such that $Q_{0}+Q_{1}=\Delta P/M$, the current can be made to vanish without invoking any external bias to counteract the density gradient. In large systems this leads to a uniform density profile and almost vanishing magnetisation in the bulk and the inhomogenities remain only in the boundary layers. Imposing an additional condition of zero density gradient ($\Delta P=0$) mimics the steady state result for reflecting boundaries [13].
  • V.

### v4

@@ Abstract
We discuss analytical results for a run-and-tumble particle (RTP) in one dimension in presence of boundary reservoirs. It exhibits ‘kinetic boundary layers’, nonmonotonous distribution, current without density gradient, diffusion facilitated current reversal and optimisation on tuning dynamical parameters, and a new transport effect in the steady state. The spatial and internal degrees of freedom together possess a symmetry, using which we find the eigenspectrum for large systems. The eigenvalues are arranged in two bands which can mix in certain conditions resulting in a crossover in the relaxation. The late time distribution for large systems is obtained analytically; it retains a strong and often dominant ‘active’ contribution in the bulk rendering an effective passive-like description inadequate. A nontrivial ‘Milne length’ also emerges in the dynamics. Finally, a novel universality is proposed in the absorbing boundary problem for dynamics with short-range colored noise. Active processes driven by active reservoirs may thus provide a common physical ground for diverse and new nonequilibrium phenomena.
[…]
@@ 1 Introduction
Systems with self-propelled or active particles became a paradigm of nonequilibrium processes. Active agents consume energy from environment and convert it into mechanical work to propel themselves. Constant energy consumption drives active systems far from equilibrium [1, 2, 3]. There are numerous real life examples of active systems across scales including bacterial colony, school of fish, flock of birds, Janus particles[4, 5], vibrated granular rods and disks [6, 7] and many more. Such systems exhibit a plethora of nontrivial collective phenomena, e.g. pattern formation, motility-induced phase separation and clustering [2, 3, 8, 9, 10], giant number fluctuations [11], Casimir effect [12]. They reveal rich nonequilibrium properties even at the single particle level: examples include non-Boltzmann steady-state, accumulation near boundary, shape transition in probability distribution in a trapping potential, a condensation-like transition in the free motion with random activity [13, 14, 15, 16, 17, 18]. Active motion is modelled with a Langevin-like equation with an ‘active noise’, which is typically coloured and violates the fluctuation-response relation, leading to such novel behaviour. Three models of scalar active particles are widely studied in literature: run-and-tumble particles (RTP), active Brownian particles (ABP) and active-Ornstein-Uhlenbeck particles (AOUP).
@@
Active particles show interesting behaviour in the presence of drives, thermal noise, obstacles and different boundary conditions. For example, RTPs subject to spatially periodic drive go through a transition from non-ergodic trapped states to moving states [19]. In presence of shear forces in two dimension, underdamped ABPs develop a boundary layer and flow reversal occurs near the wall [20]. Confining boundaries are known to have strong effects on active systems, e.g. boundary accumulation and boundary layer formation [13, 21], long-range effects and system size dependent behaviour [22, 23, 24, 25, 26, 27], etc. Interestingly, the nature of the boundary layer and the pressure exerted by the particles on a wall depend on the shape of the wall [21] as well as on the interactions [22]. Thermal fluctuations, ubiquitous but usually neglected, can have nontrivial effects on active motion e.g. formation of bound states [28], current modulation on a ratchet [29], and triggering phase separation [30, 31].
@@
In this work we focus on boundary-driven processes. Systems connected to particle or energy reservoirs have been studied for centuries, but there is a renewed interest in these processes particularly in the context of low dimensional transport. The canonical picture is given by Brownian motion in a finite system connected to particle baths: the steady state density profile is a linear interpolation of the densities at the boundaries [32] while the current is proportional to the density gradient as given by the Fick’s law. However in the presence of interaction, bulk drive, or more than one conserved quantities etc. richer and often nonintuitive behaviour is observed. Models of boundary driven interacting diffusive systems exhibit long-range correlation [33, 34, 35]; introduction of a bulk drive gives rise to different phases [36, 37]. The thermal transport become anomalous and often violate Fourier’s law in one dimensional interacting anharmonic chains [38, 39, 40, 41]. Nonlinear bulk density profile and nongradient steady state current emerges in simple noninteracting cases where the particle dynamics is governed by strongly correlated noise, for instance Levy walk connected to particle reservoirs [42, 40]. In contrast to the above cases where density or temperature gradient drives the current, a finite thermal conduction without temperature gradient was demonstrated for nonlinearly coupled passive systems driven by non-Gaussian noises [43]. New effects induced by ‘active heat baths’ in a passive harmonic (linearly coupled) chain, e.g. current reversal in the energy transport and boundary features in the steady state, have also been reported recently [44, 45]. In contrast, results for active systems in presence of boundary reservoirs are rare[46], and to our knowledge the effects of active reservoirs are not yet discussed. It is natural to ask what are the steady state, transport properties, dynamical fluctuations and relaxation behaviour of different boundary driven active systems and how do they differ across models and from their passive counterpart. To explore the nonequilibrium physics in this class of processes, we consider an elementary model amenable to simple analysis and exact results.
@@
In this paper we study non-interacting free RTPs in presence of thermal noise in a one-dimensional channel connected to particle reservoirs at both ends maintaining fixed particle densities and magnetisation  [^1] In $1$-dimension each RTP tumbles between two possible internal orientations; the magnetisation is defined as the difference of densities of the differently oriented particles.. The results are surprisingly rich and intriguing. The solution for steady state and large time distribution show the existence of ‘kinetic boundary layers’ [47] characterised by exponentially decaying components in the density profile near the reservoirs. These are distinct from the usual boundary layers reported earlier for confining walls or obstacles in that, here these occur in the presence of particle fluxes across the boundaries and can correspond to both accumulation and depletion. For zero boundary magnetisation, the steady state current and bulk density profile show a passive-like behaviour while the effect of activity, which is notable for large persistence, becomes prominent at the boundary layers. In the bulk, the signature of activity appears through a constant magnetisation. Remarkably, in the presence of nonzero boundary magnetisations, the density profile can become nonmonotonic, there is current without density gradient, and in particular a diffusion-facilitated current reversal and current maximisation upon tuning the particle parameters can take place. In this connection a new transport behaviour and related coefficient are identified.
@@
The time-dependent solution to the boundary value problem is identical to that with absorbing boundaries, which is relatively less discussed for RTPs [48, 49, 50, 51]. We observe that the boundary condition endows the system with a ‘reflection symmetry’. Using this, the eigenspectrum and time-dependent solutions are obtained analytically for large system-sizes. The eigenvalues of the time evolution operator is arranged in two distinct bands separated by the tumble rate. The relaxation is diffusive for large system size $L$; as $L$ is reduced the bands start overlapping, and for small enough systems the relaxation rate crosses over to an $L$-independent value. Such crossovers had been found earlier for one and two particle systems on a ring[52, 28]. We observe that in the eigenstate the ‘active’ contribution appears only as $O(L^{-1})$ correction to the leading ‘passive-like’ term for large $L$; but, contrary to expectations, at large times both the active and passive contributions occur at the same order in the bulk. For higher persistence, the large distance late time distribution and related properties are in fact dominated by the active contribution. Interestingly, although the distribution satisfies the absorbing condition at the boundaries in the presence of thermal noise, a ‘Milne length’ [53, 54, 55] is still identified. This length quantifies the apparent shift in the position of an absorbing boundary, when seen from afar, from its true position. A nonzero Milne length has also been found recently in the context of survival probability of random flights [56] and athermal RTPs [50]. In the present case it depends nontrivially on the diffusivity and particle parameters and plays a significant role both in the steady state and dynamics. A kinetic boundary layer appears in the time-dependent solution as well and this is argued to be a consequence of thermal diffusion. We further argue that this is related to a net imbalance of the differently oriented particles in the system, which goes away as soon as the diffusivity vanishes. As a whole, the interplay of thermal noise and active motion brings important ramifications in the dynamics in the presence of boundary driving.
@@
It turns out that many of the said features obtained for RTPs also appear in underdamped passive particle, athermal AOUP, and ABP, and are possibly generic to dynamics with coloured noise. In the absorbing boundary problem the emergence of kinetic boundary layers and Milne length were known long back for the underdamped passive motion [55, 57, 58, 59]. The large time density profile in that case [60] is in fact quantitatively similar to that obtained here for RTPs. It prompts us to propose that the late time distribution away from the absorbing wall is independent of the realisation of the noise, provided that the noise correlation is short ranged.
@@
Rest of the paper is organized as follows. In Section 2, we describe the model system of a boundary-driven run-and-tumble particle. The steady-state and the time-dependent properties are discussed in Sections  3 and  4, respectively. A universality in the large time distribution is proposed in Section 5. In Section 6 we summarise our perspective and conclude.
@@ 2 Model: boundary driven RTPs
We consider a run-and-tumble particle (RTP) moving in a finite one-dimensional box of length $L$ subject to a translational noise. The box is connected to particle reservoirs at the two ends. We assume that the microscopic details of the interaction of the reservoir with the system at the boundary is not important except that it provides a fixed boundary condition. The equation of motion of the particle is,
@@
where $v$ is the self-propulsion speed, $\sigma(t)$ is the intrinsic orientation (spin) at time $t$ characterized by a telegraphic noise that takes values $\pm 1$ and it flips between the two states with rate $\omega$, and $\eta(t)$ is the Gaussian white noise with $\langle\eta(t)\rangle=0,~\langle\eta(t)\eta(t^{\prime})\rangle=2D\delta(t-t^{\prime}),~D$ being the diffusion constant.
Figure 1: Schematic of the boundary driven active system. An RTP moves on a one-dimensional line bounded between $x=0$ and $L$. The system is connected to particle reservoirs at both the ends maintaining fixed particle densities $P_{0}$ and $P_{1}$ and magnetisation $Q_{0}$ and $Q_{1}$ at x=0,Lx=0,\,L respectively.
The run and tumble motion described in terms of $(x(t),\sigma(t))$ is a Markov process. The state of the particle at a time $t$ is given by, $|P(x,t)\rangle=\begin{bmatrix}P_{+}(x,t)\\ P_{-}(x,t)\end{bmatrix}$, where $P_{+}(x,t)$ and $P_{-}(x,t)$ are the probabilities to find a particle at $x$ at time $t$ with instantaneous spin +1+1 and 1-1 respectively. $P_{+}$ and $P_{-}$ satisfy the Master equations,
@@
$P_{0}^{\pm}$ and $P_{1}^{\pm}$ are the probability densities of particles with spin ++ and - at the two boundaries $x=0$ and $x=L$ respectively. The total probability density is $P(x,t)=P_{+}(x,t)+P_{-}(x,t)$. We also define a quantity $Q(x,t):=\langle\sigma\rangle(x,t)=P_{+}(x,t)-P_{-}(x,t)$ which is analogous to magnetization. In terms of $P$ and $Q$ the boundary conditions are rewritten as,
@@ 3 Steady state distribution
$D\partial_{x}^{2}P-v\partial_{x}Q=0$ (5)
$D\partial_{x}^{2}Q-v\partial_{x}P-2\omega Q=0.$ (6)
The solution to Eqs. (5)-(6) for the density and magnetisation profiles and the current $J_{0}=-D\partial_{x}P+v\,Q$ subject to the boundary conditions Eq. (4) are,
@@
where, $\Delta P=P_{1}-P_{0},~\Delta Q=Q_{1}-Q_{0}$, $D_{e}=D+\frac{v^{2}}{2\omega}$, $\mu=\frac{\sqrt{2\omega D_{e}}}{D}$, $M=\frac{v}{\mu D}\frac{1-e^{-\mu L}}{1+e^{-\mu L}}$, and $B_{0}=\frac{\Delta P-M(Q_{0}+Q_{1})}{2(M+L\frac{\omega}{v})}+Q_{0}-\Delta Q\,\frac{e^{-\mu L}}{1-e^{-\mu L}}\,$. The terms proportional to $e^{\pm\,\mu x}$ represent the ‘kinetic boundary layers’ of width $\mu^{-1}$. Here $D_{e}$ is the effective diffusivity for free RTP motion; $M$ is related to a Péclet number $Pe=\frac{v}{\sqrt{2\,\omega D}}$ as $M=\frac{Pe}{\sqrt{Pe^{2}\,+\,1}}$ for $L\gg\mu^{-1}\,$; later we shall see that $M$ is the ratio of Milne length and persistence length $l_{p}=\frac{v}{2\,\omega}$; and $B_{0}$ is an integration constant related to the boundary layer amplitude. In the $D=0$ case $\mu$ diverges and the exponential terms vanish. This gives rise to discontinuities in the density and magnetisation profiles at the boundaries which is smoothened to boundary layers for $D>0$. Below we discuss a few implications of the Eqs. (7)-(9).
I. New features in the distribution:
The steady state density and magnetisation profiles are shown in Figure 2222The probability of finding the particle in the system is, $\int_{0}^{L}P(x)\,dx=\frac{P_{0}+P_{1}}{2}L$, which is identical to that of a passive particle. This constrains $P_{0}$ and $P_{1}$ to $O(L^{-1})$ or less since the total probability cannot exceed $1$. We choose the boundary values accordingly.; we can identify the boundary layers as given by the $e^{\pm\,\mu\,x}$ terms in Eqs. (7)-(8). For $L\gg\mu^{-1}$, the magnetisation in the bulk (i.e. $\mu^{-1}\ll x\ll L$) assumes a constant value $Q_{b}=\frac{v}{2\omega D_{e}}J_{0}$. In this region the density is linear in $x$ and the density gradient which equals $\frac{-J_{0}}{D_{e}}$ is proportional to $(L+2\,l_{M})^{-1}$, where $l_{M}=\lim_{L\gg\mu^{-1}}M\,\frac{v}{2\,\omega}=\frac{v^{2}}{2\omega\,\sqrt{2\omega D_{e}}}$ is the Milne length as shown later. This increased effective system length suggests that the behaviour in the bulk pertains to boundary conditions at a distance $l_{M}$ beyond the actual boundaries. This naive picture based on the usual notion of the Milne length however holds only for $Q_{0}=Q_{1}=0$ (section 3.1): in this case the linear bulk profile can be extrapolated to $x=-l_{M}$ to the left and $x=L+l_{M}$ to the right to satisfy the boundary values, implying that the density far from the boundaries is mimicked by that of a passive dynamics in a slightly larger system. For nonzero boundary magnetisations this picture no longer works.
A connection of the Milne length to the boundary layers can be observed. Note that the density profile is linear for $D=0$ and is discontinuous at the boundaries. For $D>0$ the linear bulk density, when extrapolated to the boundaries, also deviates from the boundary condition; this deviation is precisely compensated by the kinetic boundary layers with strength proportional to $l_{M}$ representing particle accumulation at one end and depletion at the other.
Finally, the density profile may become nonmonotonic near the boundaries. In fact, for given boundary conditions the distribution can occur in two distinct categories for different parameter values, one with nonmonotonicities (Figure 2 left panel) and another being monotonous (Figure 2 right panel). Crucially, with parameters for which the current is co-directed with the boundary magnetisation(s) give rise to the nonmonotonous profile [^3] It can also be shown that the extrema occurs $O(\ln L)$ away from the boundaries.. Below we discuss the properties of the steady state current in the large $L$ limit.
Figure 2: Left panel: Density (solid line) and magnetisation (dashed line) profiles in the steady state of the boundary driven RTP. Here $v=1.0,D=1.0,w=0.3,L=10.0$ and $P_{0}=0.075,P_{1}=0.125,Q_{0}=Q_{1}=0.05$. The distribution is nonmonotonous and the slope in the bulk is negative, that corresponds to a current from the left to the right. Right panel: Density and magnetisation plots with the same boundary conditions and system size, but now with particle parameters $v=1.0,D=1.0,w=5.0$. The density profile is monotonous in this case and the slope in the bulk is positive corresponding to a current in the reverse direction. The boundary layers can be identified in both the cases.
II. Magnetisation induced current and a new transport coefficient:
The current in Eq. (9) has two components, one is due to the global density gradient consistent with the Fick’s law, and the other is a new contribution due to the boundary magnetisations. Consider $\Delta P=0$: The current is still nonzero, $J_{0}=\frac{D_{e}}{L+2\,l_{M}}\,M(Q_{0}+Q_{1})$. This prima facie suggests that the boundary magnetisations act as drives giving rise to a current in the system. This way of understanding obviously works for $\omega=0$ (i.e. the driven passive case) for which both $P_{+}$ and $P_{-}$ are locally conserved. In that case the density is uniform except at the boundary layers, and the values of the current and bulk magnetisation are simply $v(Q_{0}+Q_{1})/2$ and $(Q_{0}+Q_{1})/2$ respectively. The active ($\omega>0$) scenario on the other hand is very different. We might expect that the effect of boundary magnetisation, that induces a biased motion $u=v\,\frac{Q_{0}}{P_{0}}$ locally (at the left boundary), would disappear beyond a length because of the tumble dynamics and overdamped situation. And indeed the influence is much reduced, but it does not vanish. Instead a residual bulk magnetisation and current both diminished to $O(L^{-1})$ compared to those for $\omega=0$ persist. This is nonobvious from a microscopic point of view  [^4] Consider diffusion on a channel of length $L$ connected at both ends to reservoirs of density $P_{0}$. If the particle is driven with a bias velocity $u$ within a length $\ell\ll L,D/u$ near the boundaries, a current $j_{\ell}=u\,P_{0}\,\ell/L$ is induced which vanishes as $\ell\rightarrow 0$ for finite $u$. Only at $u\to\infty$ a finite current can be sustained, e.g. RTP in the white noise limit: $\lim_{v\to\infty,\omega\to\infty}\,\frac{v^{2}}{2\,\omega}=D_{a}\,$. In this limit, $u=v\,Q_{0}/P_{0}\to\infty$, and $J_{0}$ is finite, although $Q_{b}\sim\omega^{-1/2}\to 0\,$..
The competing components give rise to an interesting transport property. With suitable boundary magnetisations such that $Q_{0}+Q_{1}=\Delta P/M$, the current vanishes. Putting it in reverse, in a setting with zero current, e.g. in a closed 1D box such that particles cannot escape, maintaining a nonzero net boundary magnetisation induces a global density difference, $\Delta P=M(Q_{0}+Q_{1})$. This phenomenon to our knowledge classifies a new transport effect with coefficient $M$ and does not have a thermodynamic analogue  [^5] A loose resemblance to a Seebeck-like effect can be noted with boundary magnetisations in place of temperatures. If the particles are charged, a voltage difference would also emerge.. In particular, in the zero-current condition both magnetisation and local density gradient vanish in the bulk, and the effect is maintained by forces confined near the boundaries only. This is a global effect, entirely different from usual near equilibrium current balance phenomena. In the latter the thermodynamic forces (gradients) themselves remain nonzero while the respective mass currents balance each other at all points.
To put this further in context, some previous works on heat transport in interacting active systems, e.g. an active chain connected to thermal baths[46] and a passive chain connected to active heat reservoirs[44], also reported competing contributions in the heat flux, which allow nontrivial current balance on tuning certain parameters. Here we argue that this is not only generic for active systems at a more elementary level but can also be leveraged to extract additional transport behaviour  [^6] In the spirit of a thermo-couple, one can make an ‘active-couple’ by configuring a closed circuit of two channels and generate particle current with nonzero net magnetisation at the junctions..
III. Current reversal and optimisation: